OpenStudy (lukecrayonz):

When an object is dropped, the distance it falls in t seconds, assuming that air resist. is negligible is given by s(t)=17t^2 where s(t) is in feet. Suppose a medic's reflex hammer is dropped from a helicopter. Find (a) how far the hammer falls in 4 sec, (b) how fast the hammer is traveling 4 sec after being dropped and (c) the hammer's acceleration after it has been falling for 4 sec

OpenStudy (lukecrayonz):

Deriv og s(t)=17t^2 is 34t

OpenStudy (lukecrayonz):

of*

OpenStudy (lukecrayonz):

@Mertsj

OpenStudy (lukecrayonz):

is it a: 272 b: 272 c: 136

OpenStudy (lukecrayonz):

@campbell_st

OpenStudy (lukecrayonz):

Or am I mixing up the equations for which is which and the answers are 136 136 and 272?

OpenStudy (campbell_st):

ok... for (a) substitute t = 4 and evaluate (b) you need the rate of change in s with respect to t so find the 1st derivative then substitute t = 4 (c) acceleration is rate of change in velocity with respect to time. So find the 2nd derivative.

OpenStudy (lukecrayonz):

34

OpenStudy (lukecrayonz):

is the second deriv

OpenStudy (campbell_st):

thats part (c) but make sure you include units... all you need is (a) and (b)

OpenStudy (lukecrayonz):

so (a)=272 (b)=136 (c)=34

OpenStudy (campbell_st):

seems to make sense.. but make sure you have the units